Brief Exercise 9-12 3Your answer is incorrect. Try again. Pina Inc. had beginning inventory of $12.505 at cost and $20,500 at retail. Net purchases were $115,121 at cost and $159600 at retail. Net markips were S 10,200, markdowns were $7,200, and sales revenue was $146,600. Compute ending inventory at cost using the LIFO retail method. (Round ratios for computational purposes to decimal place, e.g. 78.7% and final answer to 0 decimal places, e.g. 28,987.) -25842] Ending inventory using LIFO retail method LINK 10 TE
Answer #1
LIFO Retail Method
Cost Retail
Beginning Inventory 12505 20500
Add: Net purchases 115121 159600
Add: Net Markups 10200
Less: Net Markdowns (7200)

Total

(excluding beginning inventory)

115121 162600
Ratio of cost of retail= 115121/162600*100= 71%

Total

(including beginning inventory)

127626 183100
Less: Net Sales (146600)
Ending inventory at retail 36500

To calculate the ending inventory at cost

Subract the beginning retail inventory from the ending retail inventory which will give you the ending inventory layer

= ending inventory at retail - beginning inventory at retail

= 36500 - 20500

=16000

Ending Inventory layers at retail pricess Cost of retail % End inventory LIFO cost

Retail prices

36500

Beginning : 20500

61%

(12505/20500*100)

20500 *61%= 12505

Ending : 16000

(36500-20500)

71%

115121/162600*100)

16000 *71%= 11360
Total: 23865